# Re-Envisioning Command and Control

### v1

@@ Abstract
Scalable Interactive Machine Learning, Artificial Intelligence, Human-AI Teaming, Future Command and Control, Human-Machine Partnerships
@@ I Introduction
A foundational challenge to re-envisioning C2 systems consists of effectively adapting a fundamentally human endeavor to include human-machine partnerships. Many modern C2 approaches arose from the era when armies outgrew their communication capabilities, prompting reformers, including 19th-century Prussian field marshal Helmuth von Moltke, to create training and operational techniques, tactics, and procedures (TTPs) that allowed for the army hierarchy to perform C2 on the principles of commander’s intent [6]. We propose here that many of the TTPs in practice today have evolved human-centric features (such as repetition, multi-form communication, group ideation, and reformulating concepts) to ensure commanders and staff under operational stressors and throughout the hierarchy have sufficient situational understanding and trust to effectively maintain unity of effort. The challenge arises in that attempts to integrate advanced technologies often focus on replacing the very human-centric features that are critical to building unity of effort within the command.
Herein, we envision a future in which the U.S. Army embraces the respective strengths and requirements of human and emerging machine intelligence to achieve enhanced C2 that outperforms decisions made by either humans or artificial intelligence (AI) alone. We will discuss the assumptions that frame the vision and then lay out a prospective C2 system capability in terms of three operational impacts: streamlining the C2 operations process [^3] The C2 operations process is the Army’s framework for organizing and putting C2 into action, and consists of the four major C2 activities performed during operations: planning, preparing, executing, and continuously assessing [7]., maintaining unity of effort, and developing adaptive collective knowledge systems.
@@ II Framing the Vision
To achieve decision dominance under the conditions that these assumptions present, most visions of future C2 systems point to the need to integration both human and machine intelligence. However, the nature of the human-machine integration that frames these concepts can be wildly different as discussed below (see Metcalfe, Perelman et al. [10] for a detailed discussion of human-centric versus machine-centric integration). At one extreme, human-centric visions focus on aiding humans to more effectively perform the fundamental C2 tasks (e.g., technologies ranging from human performance-enhancing techniques to individualized, commander intelligent decision support tools). At the other extreme, “machine-centric” visions focus on leveraging advanced intelligent systems to replace Soldiers in future C2 systems (e.g., replacements may include many of the command staff that perform the fundamental C2 tasks that underlie critical decisions). We forego both extremes and propose that the complexity, dynamics, and challenges of the future operating environment will force the effective integration of significant human resources within C2, while continuous technological advancements will force fundamental shifts in the roles and actions of humans in future C2. We frame the problem as focusing on creating human-machine partnerships [10] capable of adapting to battlefield changes while inherently enabling unity of effort across large heterogeneous groups of humans and machines.
@@ III Re-envisioning C2
Recent breakthroughs in interactive machine learning (ML) and hybrid human-machine intelligence have demonstrated that humans and AI working together can outperform either humans or AI alone in a range of domains [11, 12, 13, 14, 15]. There is, however, no one-size-fits-all approach to human-machine integration. Metcalfe, Perelman et al. [10] introduce a landscape of human-AI partnership, which posits that the nature of human-AI interaction depends on task complexity, decision timescales, and information certainty. While some simpler sub-tasks or functions may best be solved by technologies alone, many of the fundamental tasks that underlie effective future C2 fall within the complex landscape requiring effective human-machine partnership. Furthermore, over time, more and more sub-tasks and functions are expected to be solved via technology, which in turn will shift the nature of the interactions on the complex landscape. These factors make it difficult or even inappropriate to specify a type of future C2 human-machine partnership. To overcome this challenge, we take a multi-pronged approach. Below, we discuss how future C2 systems can potentially impact three specific aspects affecting the future C2 operations process. We augment each of these discussions with a future Soldier “account” aimed at inspiring imagination. Finally, we augment this discussion with two companion papers also submitted to ICMCIS 2024. The first is a research paper describing a large language model-driven COA generation capability (see Goecks and Waytowich) that illustrates the near-term capability at the heart of the examples herein, and the second outlines a broad research program illustrating the types of science and technology that potentially will enable the discussed technologies herein (see Madison, Novoseller, Goecks, et al.).
@@
My first military occupational specialty in the Army was data wrangling. I examined battlefield data, e.g., images and audio, and compared them to the machine’s interpretations. At first it was simple stuff like targets, but over time the machines got better, and later I looked at whole phases of an operation for concepts like danger. To this day, I still data wrangle for short “sprints” when major events occur. For my next deployment, I was assigned to an intelligence functional cell. In this role, I would analyze potential threat and friendly COAs using specialty AI tools that would advise me on potential intelligence-specific risks and benefits. The tools also allowed me to ask questions to help me understand the rationale behind the risk and benefit analyses as well as allowing me to compare alternatives. This was a pretty intense job, since new potential COAs were regularly being generated as new information arose. Our goal was to be familiar enough with viable possibilities that we could clearly articulate an intelligence recommendation on any COA that we were asked about. Following intelligence, I spent time in fires and sustainment cell analysis teams. Through these three roles, I felt I had a pretty good idea of how to analyze a COA. I am now on a COA development team within the Future Operations integrating cell, and it’s a way more dynamic process! The machines are great at generating and adjusting COAs as new information arrives and is verified. I have had several roles on this team. I first focused on ensuring the COAs were aligned to mission intent and parameters. In another role, I was part of a team that would identify which COA options are potentially viable, and I liaised with the analysis teams to evaluate these options and generate a cohesive assessment. Right now, I perform COA adjustments. In this role, I have many tools to create options including: speaking to a plan generator, sketching on a board, moving assets in a virtual sandbox, and/or just highlighting part of a plan and saying “good,” “bad,” “fix that.” In each case, the machines immediately adjust the COA and provide feedback on risks and benefits, and together we quickly iterate until we reach a solution.
III-A Streamlining C2 Operations Process
@@
We anticipate that human-AI interaction will streamline C2 operations process (i.e., combining multiple steps in a process or executing processes simultaneously) in contrast to the current C2 approach, which is more linear and consists of many sequential steps [8, 1]. One of the primary outcomes of streamlining C2 is to dramatically reduce the timelines of deliberate planning by enabling simultaneous COA development and analysis, as well as optimizing elements of control through rapidly generating and selecting branches and sequels [^5] Branches and sequels adjust plans and orders beyond the initial stages of an operation. Branches are “contingency options built into the base plan used for changing the mission, orientation, or direction of movement of a force…based on anticipated events, opportunities, or disruptions caused by enemy actions and reactions” [8]. A sequel is “the subsequent operation or phase based on the possible outcomes of the current operation or phase” [8], for instance a counteroffensive following a defense..
@@
At its core, we envision an interactive technological capability, or Intelligent Course of Action Suite (iCOAs), that can almost instantly integrate available information into a high-resolution Course of Action (COA) that contains much more detailed information compared to current COAs. Such a capability includes multiple intelligent technologies and processes that simultaneously create, analyze, and compare plans along an array of mission and operational variables and perspectives critical to operations (e.g., fires, maneuver, and sustainment warfighting functions; moral, ethical, sociocultural, and political implications; adversarial, civilian, and friendly actions; etc.). Further, it is expected that the iCOAs runs in real-time to process data from battlefield sensors and sources to maintain up-to-date situational awareness and generate branches and sequels as needed. Early instantiations of the suite are likely to heavily rely on humans to refine, analyze, and compare the plans; however, as the technologies evolve, we foresee a highly integrated human-machine partnership with both humans and machines filling multiple roles in C2 (see SSG Hill).
@@
This vision builds on ideas proposed by Farmer [9], which applies calculations typically performed during COA analysis within the COA development step to rapidly compare and optimize over many COA alternatives in an example scenario. In contrast, we anticipate an iCOAs that not only leverages the search capabilities of AI, but also draws on the combined strength of humans and AI to maximize creativity, exploration, military expertise, and common sense across the C2 operations process. The envisioned streamlined iCOAs will present and explain proposed COAs to personnel in an intuitive format, for instance via modalities such as language, annotated maps, or trees of possible outcomes, while providing standard COA analysis outputs such as synchronization matrices [16] and metrics such as success probabilities and uncertainty estimates that account for environmental variables and potential threat COAs. C2 personnel will input data to the system as it becomes available, analyze presented COA alternatives and associated data, interactively adapt AI-generated COAs based on domain expertise, common sense, and changing situational needs, and make final COA selections.
@@
C2 personnel will dynamically interact with the AI system to adapt proposed COAs in a variety of ways depending on situational needs. C2 personnel could describe objectives (e.g., “Take control of the enemy asset”), issue corrections as needed to adapt and fine-tune proposed plans (“approach from the opposite direction to avoid being seen”), and specify constraints that prohibit the AI from choosing undesirable actions (e.g., “don’t approach too close to the mountain”). When unanticipated changes occur, personnel could relay feedback to the AI system to indicate how to adapt the plan to emergent risks and opportunities (e.g., “the enemy has weaponized trees, so avoid trees when possible”). C2 personnel could adjust COA selection criteria to trade-off between multiple objectives, for instance upweighting the importance of controlling an enemy asset or penalizing actions that exacerbate danger to personnel or equipment, and could select among multiple COA alternatives to teach the system which COA features are preferable. Ultimately, a COA would be selected once the commander is satisfied with the presented alternatives; this COA could continue to be updated at multiple decision points and as battlefield conditions evolve.
@@
The envisioned capability could allow for compression of battle rhythms [^6] Battle rhythms are a deliberate daily cycle of command, staff, and unit activities established by commanders, consisting of a series of meetings, briefings, and other activities synchronized by time and purpose [1]. though more sophisticated data processing pipelines, leading to faster actionable intelligence for more efficient decision-making. The high-resolution nature of iCOAs would allow for faster adoption of COAs down the chain of command. Systems that enable such streamlining would accelerate C2 operations and potentially reduce the physical footprint of the Army on the battlefield by reducing necessary personnel and the required resources for logistics and sustainment. In addition to deliberate planning, the real-time nature of the capabilities allows for the same capabilities to be used during both initial planning and dynamic operations. Running continuously, the iCOAs would update predicted outcomes of COAs based on battlefield progress and generate potential adjustments as unanticipated events and information emerge. These capabilities would enable decision-making at fast enough speeds to enable forces to exploit or create brief windows of opportunity and quickly adapt to dynamic battlefield conditions to gain decision advantage.
@@ III-B Maintaining Unity of Effort
“Alert, alert, communications are actively being jammed!” We knew they would jam us, but my iCOAs-S is indicating it occurring much earlier than expected. We had not anticipated adversary activity until we had at least reached phase line PL Coppermine. Suddenly, a muted boom, and we can see the skyline light up with rocket trails and flashes of explosions—it looks like they are trying to block the entire movement corridor. Using on-board sensors, our COP [^7] Common Operating Picture is indicating that the attack is targeting an area approximately six klicks to the north-north-east and covers an area at least 20 klicks wide. It is also indicating that there is potentially active fire over 30 klicks to the east, south-east, and south. I focus on the iCOAs-S: it indicates that our current COA is suboptimal to meet our mission objectives and has three alternatives prioritized by expected success rate. The first two COAs have a much higher success rate, and so I focus on analyzing those. While I am doing my analysis, I ask iCOAs-S what’s the primary difference between the COAs. It responds that COA Alpha is a coordinated northern-focused attack that is more conservative, will only partially attain mission goals, and has a slightly higher expected casualty rate. It is highly likely that all units in this operation will have COA Alpha as a top alternative, and within the next 5-10 minutes via on-board sensors, units will be able to confirm high unity of effort. COA Bravo is a multi-pronged attack requiring northern, eastern, and southern movements. If successful, it will fully attain mission goals and has a slightly lower expected casualty rate. However, it is unlikely that the 773th Battalion operating far to the west would have COA Bravo as an alternative given the predicted range of the jamming, terrain, weather, and the capabilities of their on-board sensory. I ask iCOAs-S for a plan to get the critical information to the 773th Battalion. “Update” flashes on the screen, on-board sensors indicate that Captain Nacua over in Zulu Company already taken action to transmit emergency information to the 773th. I don’t fully understand how the technology works, but it can get very small amounts of critical information through jamming at some increased risk to the unit. I can see that COA Bravo’s expected success rate is increasing rapidly. I quickly work with iCOAs-S to refine my unit’s specific COA and order our unit to starting executing Bravo. For the next 5 minutes, I monitor the unity of effort meter and the emergency communications channel, and it seems everyone is coordinated.
@@
Critically, we envision that iCOAs-S would maintain COAs across multiple hierarchical levels while integrating information from each echelon. Running locally at each echelon, this “high-fidelity” resource would be able to predict enemy and friendly behaviors in unexpected operational events across a far broader scope than is currently expected of even the best unit leaders. We propose that this enhanced predictability (tailored to the landscape of human-machine partnership) inherently allows units to maintain unity of effort, and has the potential to dramatically accelerate coordinated action and reduce communication requirements.

### v2

@@ Abstract
Interactive Machine Learning, Artificial Intelligence, Human-AI Teaming, Command and Control, Human-Machine Partnerships
@@ I Introduction
A foundational challenge to re-envisioning C2 systems consists of effectively adapting a fundamentally human endeavor [^3] Referring to the nature of the endeavor as focused on human behavior even if aided by technology. to include human-machine partnerships. Many modern C2 approaches arose from the era when armies outgrew their communication capabilities, prompting reformers, including 19th-century Prussian field marshal Helmuth von Moltke, to create training and operational techniques, tactics, and procedures (TTPs) that allowed for the army hierarchy to perform C2 on the principles of commander’s intent [6]. We propose here that many of the TTPs in practice today have evolved human-centric features (such as repetition, multi-form communication, group ideation, and reformulating concepts) to ensure commanders and staff under operational stressors and throughout the hierarchy have sufficient situational understanding and trust to effectively maintain unity of effort. The challenge arises in that attempts to integrate advanced technologies often focus on replacing the very human-centric features that are critical to building unity of effort within the command.
Herein, we envision a future in which the U.S. Army embraces the respective strengths and requirements of human and emerging machine intelligence to achieve enhanced C2 that outperforms decisions made by either humans or artificial intelligence (AI) alone. We will discuss the assumptions that frame the vision and then lay out a prospective C2 system capability in terms of three operational impacts: streamlining the C2 operations process, [^4] The C2 operations process is the Army’s framework for organizing and putting C2 into action, and consists of the four major C2 activities performed during operations: planning, preparing, executing, and continuously assessing [7]. maintaining unity of effort, and developing adaptive collective knowledge systems.
@@ II Framing the Vision
To achieve decision dominance under the conditions that these assumptions present, most visions of future C2 systems point to the need to integrate both human and machine intelligence. However, the nature of the human-machine integration that frames these concepts can be wildly different as discussed below (see Metcalfe, Perelman et al. [10] for a detailed discussion of human- versus machine-centric integration). At one extreme, human-centric visions focus on aiding humans to more effectively perform the fundamental C2 tasks (e.g., technologies ranging from human performance-enhancing techniques to individualized, intelligent commander decision support tools). At the other extreme, machine-centric visions focus on leveraging advanced intelligent systems to replace Soldiers in future C2 systems (e.g., replacements may include many of the command staff that perform the fundamental C2 tasks that underlie critical decisions). We forego both extremes and propose that the complexity, dynamics, and challenges of the future operating environment will force the effective integration of significant human resources within C2, while continuous technological advancements will force fundamental shifts in the roles and actions of humans in future C2. We frame the problem as focusing on creating human-machine partnerships [10] capable of adapting to battlefield changes while inherently enabling unity of effort across large heterogeneous groups of humans and machines.
@@ III Re-envisioning C2
Recent breakthroughs in interactive machine learning (ML) and hybrid human-machine intelligence indicate the potential for humans and AI to work together to leverage their respective strengths [11, 12, 13]. There is, however, no one-size-fits-all approach to human-machine integration. Metcalfe, Perelman et al. [10] introduce a landscape of human-AI partnership, which posits that the nature of human-AI interaction depends on task complexity, decision timescales, and information certainty. While some simpler sub-tasks or functions may best be solved by technologies alone, many of the fundamental tasks that underlie effective future C2 fall within the complex landscape requiring effective human-machine partnership. Furthermore, over time, more and more sub-tasks and functions are expected to be solved via technology, which in turn will shift the nature of the interactions on the complex landscape. These factors make it difficult or even inappropriate to specify a type of future C2 human-machine partnership. To overcome this challenge, we take a multi-pronged approach. Below, we discuss how future C2 systems can potentially impact three specific aspects affecting the future C2 operations process. We augment each of these discussions with a future Soldier “account” aimed at inspiring imagination. Finally, we augment this discussion with two companion papers [14, 15]. The first is a research paper describing a large language model-driven course of action (COA) generation capability (see Goecks and Waytowich [15]) that illustrates the near-term capability at the heart of the examples herein, and the second outlines a broad research program illustrating the types of science and technology that potentially will enable the discussed technologies herein (see Madison, Novoseller, Goecks, et al. [14]).
@@
My first military occupational specialty in the Army was data wrangling. I examined battlefield data, e.g., images and audio, and compared them to the machine’s interpretations. At first it was simple stuff like targets, but over time the machines got better, and later I looked at whole phases of an operation for concepts like danger. To this day, I still data wrangle for short “sprints” when major events occur. For my next deployment, I was assigned to an intelligence functional cell. In this role, I would analyze potential threat and friendly COAs using specialty AI tools that would advise me on potential intelligence-specific risks and benefits. The tools also allowed me to ask questions to help me understand the rationale behind the risk and benefit analyses as well as allowing me to compare alternatives. This was a pretty intense job, since new potential COAs were regularly being generated as new information arose. Our goal was to be familiar enough with viable possibilities that we could clearly articulate an intelligence recommendation on any COA that we were asked about. Following intelligence, I spent time in fires and sustainment cell analysis teams. Through these three roles, I felt I had a pretty good idea of how to analyze a COA. I am now on a COA development team within the Future Operations integrating cell, and it’s a way more dynamic process! The machines are great at generating and adjusting COAs as new information arrives and is verified. I have had several roles on this team. I first focused on ensuring the COAs were aligned to mission intent and parameters. In another role, I was part of a team that would identify which COA options are potentially viable, and I liaised with the analysis teams to evaluate these options and generate a cohesive assessment. Right now, I perform COA adjustments. In this role, I have many tools to create options, including: speaking to a plan generator, sketching on a board, moving assets in a virtual sandbox, and/or just highlighting part of a plan and saying “good,” “bad,” “fix that.” In each case, the machines immediately adjust the COA and provide feedback on risks and benefits, and together we quickly iterate until we reach a solution.
III-A Streamlining the C2 Operations Process
@@
We anticipate that human-AI interaction will streamline the C2 operations process (i.e., combining multiple steps in a process or executing processes simultaneously) in contrast to the current C2 approach, which is more linear and consists of many sequential steps [8, 1]. One of the primary outcomes of streamlining C2 is to dramatically reduce the timelines of deliberate planning by enabling simultaneous COA development and analysis, as well as optimizing elements of control through rapidly generating and selecting branches and sequels.  [^6] Branches and sequels adjust plans and orders beyond the initial stages of an operation. Branches are “contingency options built into the base plan used for changing the mission, orientation, or direction of movement of a force…based on anticipated events, opportunities, or disruptions caused by enemy actions and reactions” [8]. A sequel is “the subsequent operation or phase based on the possible outcomes of the current operation or phase” [8], for instance a counteroffensive following a defense.Such streamlining has the potential to bring greater rigor to C2 processes, e.g., increasing analytic power with fewer available personnel, which could create decision advantage when the need for hasty, short duration planning arises (e.g., Crisis-Action Planning).
@@
At its core, we envision an interactive technological capability, or Intelligent Course of Action Suite (iCOAs), that can almost instantly integrate available information into a high-resolution COA that contains much more detailed information compared to current COAs. Such a capability includes multiple intelligent technologies and processes that simultaneously create, analyze, and compare plans along an array of mission and operational variables and perspectives critical to operations (e.g., fires, maneuver, and sustainment warfighting functions; moral, ethical, sociocultural, and political implications; adversarial, civilian, and friendly actions; etc.). Further, it is expected that the iCOAs runs in real-time to process data from battlefield sensors and sources to maintain up-to-date situational awareness and generate branches and sequels as needed. Early instantiations of the suite are likely to heavily rely on humans to refine, analyze, and compare the plans; however, as the technologies evolve, we foresee a highly integrated human-machine partnership with both humans and machines filling multiple roles in C2 (see SSG Hill).
@@
This vision builds on ideas proposed by Farmer [9], which applies calculations typically performed during COA analysis within the COA development step to rapidly compare and optimize over many COA alternatives in an example scenario. In contrast, we anticipate an iCOAs that not only leverages the search capabilities of AI, but also draws on the combined strength of humans and AI to maximize creativity, exploration, military expertise, and common sense, as well as address ethical dilemmas, across the C2 operations process. The envisioned streamlined iCOAs will present and explain proposed COAs to personnel in an intuitive format, for instance via modalities such as language, annotated maps, or trees of possible outcomes, while providing standard COA analysis outputs such as synchronization matrices [16] and metrics such as success probabilities and uncertainty estimates that account for environmental variables and potential threat COAs. C2 personnel will input data to the system as it becomes available; analyze presented COA alternatives and associated data; interactively adapt AI-generated COAs based on domain expertise, common sense, and changing situational needs; and make final COA selections.
@@
C2 personnel will dynamically interact with the AI system to adapt proposed COAs in a variety of ways depending on situational needs. C2 personnel could describe objectives (e.g., “Take control of the enemy asset”), issue corrections as needed to adapt and fine-tune proposed plans (“approach from the opposite direction to avoid being seen”), and specify constraints that prohibit the AI from choosing undesirable actions (e.g., “maintain a distance between 800 to 1000 m above mountain terrain”). When unanticipated changes occur, personnel could relay feedback to the AI system to indicate how to adapt the plan to emergent risks and opportunities (e.g., “the enemy has weaponized trees, so avoid trees when possible”). C2 personnel could adjust COA selection criteria to trade-off between multiple objectives, for instance upweighting the importance of controlling an enemy asset or penalizing actions that exacerbate danger to personnel or equipment, and could give feedback on multiple COA alternatives to teach the system which COA features are preferable. It is also envisioned that C2 personnel and the iCOAs each have the capability to ask for help from the other on decisions, tasks, or challenges as needed. Ultimately, a COA would be selected once the commander is satisfied with the presented alternatives; this COA could continue to be updated at multiple decision points and as battlefield conditions evolve.
@@
The envisioned capability could allow for compression of battle rhythms [^7] Battle rhythms are a deliberate daily cycle of command, staff, and unit activities established by commanders, consisting of a series of meetings, briefings, and other activities synchronized by time and purpose [1]. though more sophisticated data processing pipelines, leading to faster actionable intelligence for more efficient decision-making. The high-resolution nature of iCOAs would allow for faster adoption of COAs down the chain of command. Systems that enable such streamlining would accelerate C2 operations and potentially change the physical footprint of the Army on the battlefield by allowing for organizational changes in the Army, changing personnel requirements, and/or altering the required resources for logistics and sustainment. In addition to deliberate planning, the real-time nature of the capabilities allows for the same capabilities to be used during both initial planning and dynamic operations. Running continuously, the iCOAs would update predicted outcomes of COAs based on battlefield progress and generate potential adjustments as unanticipated events and information emerge. These capabilities would enable decision-making at fast enough speeds to enable forces to exploit or create brief windows of opportunity and quickly adapt to dynamic battlefield conditions to gain decision advantage.
@@ III-B Maintaining Unity of Effort
“Alert, alert, communications are actively being jammed!” We knew they would jam us, but my iCOAs-S is indicating it’s occurring much earlier than expected. We had not anticipated adversary activity until we had at least reached phase line PL Coppermine. Suddenly, a muted boom, and we can see the skyline light up with rocket trails and flashes of explosions—it looks like they are trying to block the entire movement corridor. Using on-board sensors, our COP [^8] Common Operating Picture is indicating that the attack is targeting an area approximately six klicks to the north-north-east and covers an area at least 20 klicks wide. It is also indicating that there is potentially active fire over 30 klicks to the east, south-east, and south. I focus on the iCOAs-S: it indicates that our current COA is suboptimal to meet our mission objectives and has three alternatives prioritized by expected success rate. The first two COAs have a much higher success rate, and so I focus on analyzing those. While I am doing my analysis, I ask iCOAs-S what’s the primary difference between the COAs. It responds that COA Alpha is a coordinated northern-focused attack that is more conservative, will only partially attain mission goals, and has a slightly higher expected casualty rate. It is highly likely that all units in this operation will have COA Alpha as a top alternative, and within the next 5-10 minutes via on-board sensors, units will be able to confirm high unity of effort. COA Bravo is a multi-pronged attack requiring northern, eastern, and southern movements. If successful, it will fully attain mission goals and has a slightly lower expected casualty rate. However, it is unlikely that the 773th Battalion operating far to the west would have COA Bravo as an alternative given the predicted range of the jamming, terrain, weather, and the capabilities of their on-board sensory. I ask iCOAs-S for a plan to get the critical information to the 773th Battalion. “Update” flashes on the screen, on-board sensors indicate that Captain Nacua over in Zulu Company already taken action to transmit emergency information to the 773th. I don’t fully understand how the technology works, but it can get very small amounts of critical information through jamming at some increased risk to the unit. I can see that COA Bravo’s expected success rate is increasing rapidly. I quickly work with iCOAs-S to refine my unit’s specific COA and order our unit to starting executing Bravo. For the next 5 minutes, I monitor the unity of effort meter and the emergency communications channel, and it seems everyone is coordinated.
@@
Critically, we envision that iCOAs-S would maintain COAs across multiple hierarchical levels while integrating information from each echelon. Running locally at each echelon, this “high-fidelity” resource would be able to include enemy and friendly anticipated actions (including considering the biases of their respective commanders) during unanticipated operational events, but with a broader inclusion of the anticipated actions held by different perspectives at each echelon (i.e., operational levels through to tactical levels). We assume that given DDIL, locally run instances of iCOAs-S would not all have the same real-time information; however, iCOAs-S would be able to project what real-time information is likely to be known by other individual units and predict behaviors accordingly (see Major Mahomes 773th Battalion example). We propose that this enhanced predictability (tailored to the landscape of human-machine partnership) inherently allows units to maintain unity of effort, and has the potential to dramatically accelerate coordinated action and reduce communication requirements.